We are given 50ml of 5L of HCl solutions if it is diluted and made five serial dilutions and then the most diluted sample (50ml) is titrated with 1m of NaOH solution. What volume will be required to neautralize the acidic solution?


1
Expert's answer
2020-09-05T14:39:48-0400

Molarity of HCl solution not given in question

so we assume that molarity of hcl solution is 1 M

volume of solution 5 litre

we take 50 ml solution of 1M hcl and make it serial dilution

a) first dilution add 50 ml of water in it then concentration of new solution is C1V1= C2V2

1×50=C2×100

=1/2 M

b) now further dilution with 100 ml of water in it then new concentration is

=1/4

c) add 100 ml water in it then

= 1/6

d) add 100 ml water again then concentration is

=1/8

e) now further add 100 ml water then concentration of this solution is

= 1/10

This is the most diluted solution and take 50 ml of it and neutralize with 1M NaOH solution

then volume of NaOH solution is :

50×.1=1×V

V=5 ml